OpenStudy (anonymous):

Rewrite as a single logaithm ln3+lny-2lnx

OpenStudy (anonymous):

\[\ln(3)+\ln(y)-2\ln(x)\rightarrow \ln(3y)-2\ln(x)\]\[\rightarrow \ln(3y)+\ln(x ^{-2})\rightarrow \ln(3yx ^{-2})\]Cheers.